# m m OTr_ #1#3(#2#3)

### v2

@@ III.1 Wigner’s friend
The gedankenexperiment that pushes the consequences of the measurement problem to the extreme is known as Wigner’s friend paradox [30]. This features an observer, the friend, performing measurements on a quantum system – say, a spin-1/2 particle – and a “superobserver”, Wigner, who has the ability to treat the particle and the friend as a joint quantum system. This highlights the ambiguity of quantum physics when it comes to measurements: while the friend supposedly sees the system “collapse” to a single outcome, Wigner describes the evolution of the isolated joint system (friend + particle) unitarily. This paradox has received renewed attention in recent years, and each interpretation offers a different solution [31, 32].
@@
But are Wigner’s friend-like scenarios specific to quantum theory, or can they arise also in other indeterministic theories provided with a measurement problem? Recently, the authors of [33] have shown classical analogues of Wigner’s friend scenarios. They build a series of classical thought experiments that involve the perfect cloning (and subsequent destruction) of a classical observer, which leads to ambiguities in the (probabilistic) state assignment analogous to the Wigner’s friend paradox in quantum theory. However, the scenarios of Ref. [33] cannot be directly applied to show that a theory provided with fundamental indeterminacy – like the one based on FIQs, here discussed – exhibits Wigner’s friend-like scenarios. In fact, as we will argue in Sect. VI, cloning is not possible even at the classical level if one does not uphold infinite precision. This renders the examples in [33] fundamentally unattainable in this context.
@@
Yet, one can still devise scenarios that are analogous to Wigner’s friend in a classical indeterministic theory. For instance, one could postulate that isolated systems always evolve symplectically and no events of actualization of the propensities can happen in that case. Recall that in our FIQs-based indeterministic interpretation, it is the collection of propensities (which are not formally probabilities) that represent the pure state of a system, i.e., the maximal existing information at a given time [18]. This closely resembles textbook quantum mechanics which establishes that the state of an isolated system evolves unitarily. On the other hand, one can think that the actualization of propensities happens exclusively upon interaction, in the same way that it happens in standard quantum physics (in contrast to spontaneous collapse models). Given this view, imagine a classical indeterminate system, which is in a fully isolated box together with an “observer” (i.e., a system that forces the actualization of some propensities of the system). After the interaction, for this observer the system will have a greater degree of determinacy. But since the box is isolated, for an external observer who plays the role of Wigner, everything that is in the box remains equally indeterminate because no actualization can happen for him. Obviously, this example is built in full analogy with quantum theory and shows an ambiguity between two different types of dynamics both at the quantum and at the (indeterministic) classical level: unitary vs. “collapse” in quantum theory, and symplectic vs. “actualization” in classical physics.
@@
There is yet another way in which objective actualization can lead to Wigner’s friend, without invoking these constraints on the dynamics. Namely, in Ref. [34], we have argued that if one wants to have fundamental indeterminacy in a relativistic scenario, it must be the case that determinacy itself is a relative fact. This means that an event of actualization happens only within the future light-cone of that event, and remains fundamentally indeterminate for space-like separated observers. In that way, one can construct a classical Wigner’s friend scenario: an “observer” measures an indeterminate system, thereby forcing it to become more determinate, whereas for a space-like separated “observer” (Wigner), the propensities of the system do not actualize due to the relativity of indeterminacy. In this way, to two observers pertain objectively different propensities, i.e., degrees of indeterminacy, associated with the same physical interaction, similar to how in quantum Wigner’s friend scenarios, to different observers pertain two different pure states.
IV A plurality of interpretations
@@
According to Copenhagen (i), the emergence of single outcomes is due to the interaction of a classical measurement device which is not described by quantum theory. According to this interpretation, for each physical situation, there is a Heisenberg cut that separates the quantum domain from the classical one, and that it is the classical apparatus that induces the appearance of a definite measurement outcome. We have already pointed out in Ref. [11] that this can be regarded as an instance of top down-causation [35, 36], namely, where a higher level of description “forces” the actualization of the lower level. In the classical case, in full analogy with the Copenhagen interpretation of quantum mechanics, one can think of a measurement apparatus as a system whose physical state have a much smaller indeterminacy than the system to be measured. This system with a higher degree of determinacy plays the role analogous to that of the measurement apparatus in the Copenhagen interpretation, i.e., the system that exhibits classical behavior on the opposite side of the Heisenberg cut. Similarly, one can conceive that it is the interaction between the higher-level (apparatus) and the lower-level (system) that forces the reduction of indeterminacy, i.e., the actualization of the state, up to the determinacy level of the apparatus. In the Copenhagen interpretation of quantum physics, for an object to be considered a measurement apparatus (which lies outside the scope of the theory’s description), it must be, in a reasonable sense, macroscopic. In contrast, in indeterministic classical physics, the necessary condition for a measurement apparatus is a certain level of determinacy, which may be unrelated to the macroscopicity of the apparatus. However, just as in standard quantum mechanics not all macroscopic systems are necessarily treated as measurement apparatuses, in indeterministic classical physics, not all more determinate systems are necessarily measurement apparatuses. So, the identification of a physical system as a measurement apparatus remains an open problem in classical physics as much as in quantum mechanics. This view, however, does not require to uphold reductionism, because it could be that the properties of the higher level are (strongly) emergent and therefore not explicable in terms of the lower level.
@@
Both within Bohmian mechanics, and many-worlds interpretation, the measurement problem is considered to be only a pseudo-problem, namely, they reject the existence of fundamental potentialities and ascribe this apparent collapse to ignorance (although with opposite stratagems). In particular, Bohmian mechanics (iii) postulates that the quantum state is incomplete and that there are therefore “hidden variables” (the particles’ positions) which are not described by quantum theory [44]. However, if one were to know these hidden variables, quantum systems would have a unique evolution, i.e., quantum physics would be deterministic. [^8] We would like to stress that David Bohm was never a determinist and this interpretation, despite bearing his name, quite betrays his message [45]. This interpretation is simply analogous to the orthodox (deterministic) interpretation of classical physics, where the real numbers are the “hidden variables” [12, 15]. Only with this completion in terms of real numbers as hidden variables, namely with the orthodox interpretation of classical mechanics, does classical theory become deterministic.
@@
The idiosyncratic many-worlds interpretation (iv) rejects the fact that measurements yield single outcomes. Its proponents believe that there only exists a quantum state of the universe that undergoes unitary dynamics. So, every time we believe we are doing a measurement, it is “simply” the universe that is splitting into as many worlds as the outcomes of the measurements. All outcomes always occur, but in parallel worlds or branches of the universal wave function. Each different outcome is perceived in a different branch, but it is only due to a partial knowledge of the actual global state of the universe. Also classical indeterministic theory can be interpreted exactly in the same way, by assuming that the propensities never actualize towards one value (ruling out the alternative outcomes), but all obtain in different worlds and are considered, in this interpretation, equally real. The whole state of the universe would thus be a collection of all the propensities for any possible event of the universe (only outcomes with propensities equal to zero do not actualize).
@@
Finally, interpretations like QBism [46] (v) take the quantum state to be a subjective piece of information used by an agent to take rational decisions about what to “bet” on. Obviously, one can take such a subjective approach also in an indeterministic interpretation of classical physics. In fact, this simply amounts to acknowledge that there is no further underlying reality about which we assign probability. In this case one can still interpret physics in terms of FIQs, but what we have taken as propensities would be interpreted as the only available piece of information, i.e. epistemic probabilities. Therefore, what we have considered so far an objective process of actualization of propensities would be a Bayesian update of the information available to an agent.
V Nonlocality and entanglement
V.1 Nonlocality à la Eintein
@@
A first problem with the nonlocality of quantum mechanics was pointed out by Einstein, even before the complete formalization of the theory [47]. Einstein considered the wave packet of a single quantum particle moving towards a screen. While the whole wave front reaches every point of the screen at the same time, the latter plays the role of a (position) measurement device, therefore “collapsing” the wave packet to a localized region. This introduces an (in principle) instantaneous correlation between the point where the particle localizes and all the other points of the screen that all of a sudden lose the possibility of detecting the particle. Equivalently, this is what happens when a single photon passes through beamsplitter that prepares its state in a coherent superposition of the two spatial locations. But if one is to measure the position, the photon will always appear only in one branch, even if the detection processes are space-like separated [48].
@@
V.2 Entanglement in indeterministic classical physics
@@
In this section we consider a classical analogue of entanglement, namely, the possibility that indeterminate systems that have interacted in the past maintain correlations in their joint state. It can so happen that the indeterminacy of the joint system is lower that the respective local indeterminacy of each of the two systems. This is for instance what happens in the case of parametric down conversion in photons: A laser of well-determined energy pumps a nonlinear crystal to produce two photons each of which has a large indeterminacy in energy (i.e., their spectra are very broad), but the sum of their energies remains that of the pumping laser, i.e., with very small indeterminacy in energy.
As a toy example in classical indeterministic physics, consider two particles whose motion is constrained in one dimension, one located with Alice and one with Bob. However, both particles have a position that, like every other physical degree of freedom, is subjected to some ontic indeterminacy quantified, respectively, by $\Delta a_{0}$ and $\Delta b_{0}$, with $\Delta a_{0}\leq\Delta b_{0}$, where the subscript index 00 means at time $t=0$. Hence the initial state of $A$ can be represented by $\omega_{A}(t=0)=a_{0}\pm\Delta a_{0}/2$, and the one of $B$ by $\omega_{B}(t=0)=b_{0}\pm\Delta b_{0}/2$ (see Fig. 4). The reference positions $a_{0}$ and $b_{0}$ are just coordinates to label the relative position with respect to an origin, but do not have a physical relevance. Note that only their combined state is pure (i.e., it encapsulates the maximal information).
Assume now, for simplicity, that every time there is a local actualization event – either due to a measurement, or spontaneously as time passes (see section IV) – the indeterminacy of the position halves. Assume, again for simplicity, that each actualization localizes the particle in either the first or the second half of the interval $\Delta a_{i}$ (but in reality it can localize anywhere within that interval). In this way, at the next actualization event, say at time $t=1$, the state actualizes as either $\omega_{A}(t=0)\rightarrow\omega_{A}^{-}(t=1)=a-\Delta a_{0}/4\pm\Delta a_{0}/4$ or $\omega_{A}(t=0)\rightarrow\omega_{A}^{+}(t=1)=a+\Delta a_{0}/4\pm\Delta a_{0}/4$. The local ontic indeterminacy will halve as $\Delta a_{1}=\Delta a_{0}/2$. Before the actualization, i.e., at $t<1$, both $\omega_{A}^{-}(t=1)$ and $\omega_{A}^{+}(t=1)$ are possible future states – in general with different propensities – but only one among the two will actualize at $t=1$.
@@
What we have here closely resembles quantum entanglement, for it shows correlations of nondeterministic variables. However, contrarily to quantum entanglement, this classical form cannot violate a Bell inequalities due to the non-existence of incompatible variables, and thus allows for the possibility of local hidden variables [15]. Note that, like for quantum entanglement, the randomness of the local measurement outcomes – in our example, whether the indeterminacy reduces to the first or the second half of the interval – does not allow to signal. However, a local actualization event instantaneously changes the state on the other side. Therefore, we have to assume, similarly to Bohmian mechanics where the hidden variables – i.e., particles’ positions – cannot be perfectly known, that the classical indeterminate states – i.e. the complete list of propensities – can also not be perfectly known (unless one allows instantaneous signaling).
@@ VI No-cloning
It was already shown that this result is not characteristic of the quantum formalism per se, but it holds unchanged for statistical ensembles of classical systems [53, 54]. Let us recap the proof provided in Ref. [54].
@@
These considerations can be straightforwardly applied to our alternative indeterministic interpretation of classical physics based on FIQs. In fact, assume that a physical quantity is represented by a FIQ as in Eq. (2). In this way, what is regarded as a probability distribution over statistical ensemble in the orthodox view, becomes here an ontic probability representing a measure of fundamental indeterminacy. [^9] There are some mathematical subtleties for which propensities are not formally Kolmogorov probabilities (see Ref. [18]), but for an intuitive understanding this distinction does not matter here. That is, the two different distributions $P_{1}^{(s)}(\textbf{x}^{(s)})$ and $P_{2}^{(s)}(\textbf{x}^{(s)})$ are now pure states in phase space which contain finite information. Therefore, we conclude that arbitrary classical states that have ontic indeterminacy cannot be copied.
@@
However, one cannot use this classical no-cloning theorem to ensure the information-theoretic security of classical key distribution protocols. The reason is that here there is no $\hbar$ and hence no incompatible measurements, which are required to generate a secure key. Indeed, to bound the information that an adversary could gain on a variable $A$, one measures the disturbance that the adversary necessarily introduces on a conjugate variable $B$.
VII Classical vs. Quantum: which one is more indeterminate?
@@
VIII Conclusions: what is really quantum is incompatibility
So far, we have shown that a number of features and conceptual problems, which are usually attributed to quantum physics, naturally arise from ontic indeterminacy, also in the framework of classical physics. This by no means wants to show that every quantum phenomenon can be explained by an indeterministic interpretation of classical physics.

### v3

@@ 3.1 Wigner’s friend
The gedankenexperiment that pushes the consequences of the measurement problem to the extreme is known as Wigner’s friend paradox [30]. This features an observer, the friend, performing measurements on a quantum system – say, a spin-1/2 particle – and a “superobserver", Wigner, who has the ability to treat the particle and the friend as a joint quantum system. This highlights the ambiguity of quantum physics when it comes to measurements: while the friend supposedly sees the system “collapse" to a single outcome, Wigner describes the evolution of the isolated joint system (friend + particle) unitarily. This paradox has received renewed attention in recent years, and each interpretation offers a different solution [31, Frauchiger2018].
@@
But are Wigner’s friend-like scenarios specific to quantum theory, or can they arise also in other indeterministic theories provided with a measurement problem? Recently, the authors of [32] have shown classical analogues of Wigner’s friend scenarios. They build a series of classical thought experiments that involve the perfect cloning of a classical observer, which leads to ambiguities in the (probabilistic) state assignment analogous to the Wigner’s friend paradox in quantum theory. However, the scenarios of Ref. [32] cannot be directly applied to show that a theory provided with fundamental indeterminacy – like the one based on FIQs, here discussed – exhibits Wigner’s friend-like scenarios. In fact, as we will argue in Sect. 6, cloning is not possible even at the classical level if one does not uphold infinite precision. This renders the examples in [32] fundamentally unattainable in this context.
@@
Yet, one can still devise scenarios that are analogous to Wigner’s friend in a classical indeterministic theory. For instance, one could postulate that isolated systems always evolve symplectically and no events of actualization of the propensities can happen in that case. Recall that in our FIQs-based indeterministic interpretation, it is the collection of propensities (which are not probabilities in the sense that they do not satisfy all Kolmogorov axioms; see [21]) that represent the pure state of a system, i.e., the maximal existing information at a given time [21]. This closely resembles textbook quantum mechanics which establishes that the state of an isolated system evolves unitarily. On the other hand, one can think that the actualization of propensities happens exclusively upon interaction, in the same way that it happens in standard quantum physics (in contrast to spontaneous collapse models). Given this view, imagine a classical indeterminate system, which is in a fully isolated box together with an “observer" (i.e., a system that forces the actualization of some propensities of the system). After the interaction, for this observer the system will have a greater degree of determinacy. But since the box is isolated, for an external observer who plays the role of Wigner, everything that is in the box remains equally indeterminate because no actualization can happen for him. Obviously, this example is built in full analogy with quantum theory and shows an ambiguity between two different types of dynamics both at the quantum and at the (indeterministic) classical level: unitary vs. “collapse" in quantum theory, and symplectic vs. “actualization" in classical physics.
@@
Finally, we would like to point out that despite the possibility to retrieve Wigner’s friend scenarios in classical indeterministic physics, the consequences of the measurement problem in quantum mechanics are more profound, due to recent no-go theorems involving the combination of Bell nonlocality with multipartite Wigner’s friend setups [33, 34, 35]. As we will comment below (Sect. 5.2), the lack of incompatible measurements makes it impossible to have classical Bell nonlocality and therefore the aforementioned extended Wigner’s friend theorem do not seem to have a classical counterpart.
4 A plurality of interpretations
@@
According to Copenhagen (i), the emergence of single outcomes is due to the interaction of a classical measurement device which is not described by quantum theory. According to this interpretation, for each physical situation, there is a cut (sometimes referred to as Heisenberg cut) that separates the quantum domain from the classical one, and that it is the classical apparatus that induces the appearance of a definite measurement outcome. In a widespread view of this interpretation, the Heisenberg cut is not a fixed, objective division at a certain scale (of, e.g., size or energy). We have already pointed out in Ref. [13] that this can be regarded as an instance of top down-causation [36, 37], namely, where a higher level of description “forces” the actualization of the lower level. In the classical case, in full analogy with the Copenhagen interpretation of quantum mechanics, one can think of a measurement apparatus as a system whose physical state have a much smaller indeterminacy than the system to be measured. This system with a higher degree of determinacy plays the role analogous to that of the measurement apparatus in the Copenhagen interpretation, i.e., the system that exhibits classical behavior on the opposite side of the Heisenberg cut. Similarly, one can conceive that it is the interaction between the higher-level (apparatus) and the lower-level (system) that forces the reduction of indeterminacy, i.e., the actualization of the state, up to the determinacy level of the apparatus. In the Copenhagen interpretation of quantum physics, for an object to be considered a measurement apparatus (which lies outside the scope of the theory’s description), it must be, in a reasonable sense, macroscopic. In contrast, in indeterministic classical physics, the necessary condition for a measurement apparatus is a certain level of determinacy, which may be unrelated to the macroscopicity of the apparatus. However, just as in standard quantum mechanics not all macroscopic systems are necessarily treated as measurement apparatuses, in indeterministic classical physics, not all more determinate systems are necessarily measurement apparatuses. So, the identification of a physical system as a measurement apparatus remains an open problem in classical physics as much as in quantum mechanics. This view, however, does not require to uphold reductionism, because it could be that the properties of the higher level are strongly emergent [38] and therefore not explicable in terms of the lower level.
@@
Both within Bohmian mechanics, and many-worlds interpretation, the measurement problem is considered to be only a pseudo-problem, namely, they reject the existence of fundamental potentialities and ascribe this apparent collapse to ignorance (although with opposite stratagems). In particular, Bohmian mechanics (iii) postulates that the quantum state is incomplete and that there are therefore “hidden variables" (the particles’ positions) which are not described by quantum theory [45]. However, if one were to know these hidden variables, quantum systems would have a unique evolution, i.e., quantum physics would be deterministic. [^7] We would like to stress that David Bohm was never a determinist and this interpretation, despite bearing his name, quite betrays his message [46]. This interpretation is simply analogous to the orthodox (deterministic) interpretation of classical physics, where the real numbers are the “hidden variables” [14, 18]. Only with this completion in terms of real numbers as hidden variables, namely with the orthodox interpretation of classical mechanics, does classical theory become deterministic.
@@
The idiosyncratic many-worlds interpretation (iv) rejects the fact that measurements yield single outcomes. Its proponents believe that there only exists a quantum state of the universe that undergoes unitary dynamics. So, every time we believe we are doing a measurement, it is “simply" the universe that is splitting into as many worlds as the outcomes of the measurements. All outcomes always occur, but in parallel worlds or branches of the universal wave function. Each different outcome is perceived in a different branch, but it is only due to a partial knowledge of the actual global state of the universe. Also classical indeterministic theory can be interpreted exactly in the same way, by assuming that the propensities never actualize towards one value (ruling out the alternative outcomes), but all obtain in different worlds and are considered, in this interpretation, equally real. The whole state of the universe would thus be a collection of all the propensities for any possible event of the universe (only outcomes with propensities equal to zero do not actualize).
@@
Finally, interpretations like QBism [47] (v) take the quantum state to be a subjective piece of information used by an agent to take rational decisions about what to “bet" on. One can take such a subjective approach also in an indeterministic interpretation of classical physics. In fact, this simply amounts to acknowledge that there is no further underlying reality about which we assign probability. In this case one can still interpret physics in terms of FIQs, but what we have taken as propensities would be interpreted as the only available piece of information, i.e. epistemic probabilities. Therefore, what we have considered so far an objective process of actualization of propensities would be a Bayesian update of the information available to an agent.
5 Nonlocality and entanglement
5.1 Nonlocality à la Eintein
@@
A first problem with the nonlocality of quantum mechanics was pointed out by Einstein, even before the complete formalization of the theory [48]. Einstein considered the wave packet of a single quantum particle moving towards a screen. While the whole wave front reaches every point of the screen at the same time, the latter plays the role of a (position) measurement device, therefore “collapsing" the wave packet to a localized region. This introduces an (in principle) instantaneous correlation between the point where the particle localizes and all the other points of the screen that all of a sudden lose the possibility of detecting the particle. Equivalently, this is what happens when a single photon passes through beamsplitter that prepares its state in a coherent superposition of the two spatial locations. But if one is to measure the position, the photon will always appear only in one branch, even if the detection processes are space-like separated [49].
@@
5.2 Entanglement in indeterministic classical physics
@@
In this section we consider a classical analogue of entanglement. It can so happen that the indeterminacy of the joint system is lower that the respective local indeterminacy of each of the two systems. This is for instance what happens in the case of parametric down conversion in photons: A laser of well-determined energy pumps a nonlinear crystal to produce two photons each of which has a large indeterminacy in energy (i.e., their spectra are very broad), but the sum of their energies remains that of the pumping laser, i.e., with very small indeterminacy in energy.
As a toy example in classical indeterministic physics, consider two particles whose motion is constrained in one dimension, one located with Alice and one with Bob. However, both particles have a position that, like every other physical degree of freedom, is subjected to some ontic indeterminacy quantified, respectively, by $\Delta a_{0}$ and $\Delta b_{0}$, with $\Delta a_{0}\leq\Delta b_{0}$, where the subscript index 00 means at time $t=0$. Hence the initial state of $A$ can be represented by $\omega_{A}(t=0)=a_{0}\pm\Delta a_{0}/2$, and the one of $B$ by $\omega_{B}(t=0)=b_{0}\pm\Delta b_{0}/2$ (see Fig. 4). The reference positions $a_{0}$ and $b_{0}$ are just coordinates to label the relative position with respect to an origin, but do not have a physical relevance. Note that only their combined state is pure, i.e., it encapsulates the maximal existing information about the physics of the system (the information at any given time remains finite).
Assume now, for simplicity, that every time there is a local actualization event – either due to a measurement, or spontaneously as time passes (see section 4) – the indeterminacy of the position halves. Assume, again for simplicity, that each actualization localizes the particle in either the first or the second half of the interval $\Delta a_{i}$ (but in reality it can localize anywhere within that interval). In this way, at the next actualization event, say at time $t=1$, the state actualizes as either $\omega_{A}(t=0)\rightarrow\omega_{A}^{-}(t=1)=a-\Delta a_{0}/4\pm\Delta a_{0}/4$ or $\omega_{A}(t=0)\rightarrow\omega_{A}^{+}(t=1)=a+\Delta a_{0}/4\pm\Delta a_{0}/4$. The local ontic indeterminacy will halve as $\Delta a_{1}=\Delta a_{0}/2$. Before the actualization, i.e., at $t<1$, both $\omega_{A}^{-}(t=1)$ and $\omega_{A}^{+}(t=1)$ are possible future states – in general with different propensities – but only one among the two will actualize at $t=1$.
@@
What we have here closely resembles quantum entanglement, for it shows correlations of nondeterministic variables. In fact, starting from the example above, one can in general define product states in any (indeterministic) theory as states associated with two or more subsystems, such that the indeterminacy of any physical quantity is entirely determined by the indeterminacy of each individual subsystem, without imposing additional constraints on the indeterminacy of relative physical quantities. States that are not product are then entangled. However, contrarily to quantum entanglement, this classical form cannot violate a Bell inequalities due to the non-existence of incompatible quantities, and thus allows for the possibility of local hidden variables [18]. Nevertheless, while this can be regarded as a good reason to maintain the “orthodox" interpretation of classical physics, we would like to stress again that the price to pay for adding these non-contextual hidden variables is to introduce infinite information. Note that, like for quantum entanglement, the randomness of the local measurement outcomes – in our example, whether the indeterminacy reduces to the first or the second half of the interval – does not allow to signal. However, a local actualization event instantaneously changes the state on the other side. Therefore, we have to assume, similarly to Bohmian mechanics where the hidden variables – i.e., particles’ positions – cannot be perfectly known, that the classical indeterminate states – i.e. the complete list of propensities – can also not be perfectly known (unless one allows instantaneous signaling).
@@ 6 No-cloning
It was already shown that this result is not characteristic of the quantum formalism per se, but it holds unchanged for statistical ensembles of classical systems [54, daffertshofer2002classical, 55]. Let us recap the proof provided in Ref. [daffertshofer2002classical].
@@
These considerations can be straightforwardly applied to our alternative indeterministic interpretation of classical physics based on FIQs. In fact, assume that a physical quantity is represented by a FIQ as in Eq. (2). In this way, what is regarded as a probability distribution over statistical ensemble in the orthodox view, becomes here an ontic probability representing a measure of fundamental indeterminacy. [^8] There are some mathematical subtleties for which propensities are not formally Kolmogorov probabilities (see Ref. [21]), but for an intuitive understanding this distinction does not matter here. That is, the two different distributions $P_{1}^{(s)}(\textbf{x}^{(s)})$ and $P_{2}^{(s)}(\textbf{x}^{(s)})$ are now representing pure states with finite information content. Therefore, we conclude that arbitrary classical states that have ontic indeterminacy cannot be copied.
@@
However, one cannot use this classical no-cloning theorem to ensure the information-theoretic security of classical key distribution protocols. The reason is that here there are no incompatible measurements, which are required to generate a secure key. Indeed, to bound the information that an adversary could gain on a variable $A$, one measures the disturbance that the adversary necessarily introduces on a conjugate variable $B$.
7 Classical vs. Quantum: which one is more indeterminate?
@@
8 Conclusions: what is truly quantum is incompatibility
So far, we have shown that a number of features and conceptual problems, which are usually attributed to quantum physics, naturally arise from ontic indeterminacy, also in the framework of classical physics. We have provided such an explicit model of classical indeterministic physics, based on FIQs. [^9] Note that despite its intuitiveness and simplicity, the way we here introduced FIQs is strictly speking not fully satisfactory since the information content is not invariant under change of units and basic arithmetic [62]. A complete definition of FIQs is provided in the language of intutionism [63, 64, 19].
Our work is similar in spirit to recent progress aimed at showing that certain features, usually attributed to quantum theory, are in fact retrieved in toy theories that are classical (under a certain definition thereof) [2, 3, 4, 5, 6]. The fundamental difference is, however, that while in Refs. [1, 2, 3, 5] the authors consider a set of ontologically classical (i.e., fully determinate) states with epistemic constraints, for us the incompatibility of physical variables at the ontic level is what characterizes non-classicality. In contrast, in the aforementioned works, the epistemic limitations allow them to find uncertainty relations in their models. The advantage of our approach is, however, that we do not reverse engineer certain sought properties inspired by quantum theory, but we simply take the complete actual theory of classical mechanics and reinterpret it indeterministically at the ontic level to show how close to quantum theory one can go.